Phosphorus pentachloride is produced by the reaction between phosphine and chlorine gas. What mass of chlorine gas would be required to produced 0.015kg of phosphorus pentachloride?
1
Expert's answer
2019-04-01T07:48:19-0400

2P + 5Cl2 → 2PCl5

M(PCl5) = 31 + 35.5*5 = 208.5 g/mol

n(PCl5) = m/M = 15g/208.5g/mol = 0.072 mol

n(Cl2) = 0.072*2.5 = 0.18 mol

m(Cl2) = n*M = 0.18mol*71g/mol = 12.78 g